# Scaling Laws for Function Diversity and Specialization Across Socioeconomic and Biological Complex Systems

### v2

@@ Abstract
Function diversity, or the range of tasks that individuals perform, is essential for productive organizations. In the absence of overarching principles, the characteristics of function diversity are seemingly unique to each domain. Here, we introduce an empirical framework and a mathematical model for the diversification of functions in a wide range of systems, such as bacteria, federal agencies, universities, corporations, and cities. Our findings reveal that the number of functions within these entities grows sublinearly with system size, with exponents ranging from 0.35 to 0.57, confirming Heaps’ Law. In contrast, cities exhibit logarithmic growth in the occupation types. We generalize the Yule-Simon model to quantify a wide range of these empirical observations by introducing two new key attributes: a diversification parameter that characterizes the tendency for more populated functions to inhibit new function creation, and a specialization parameter that describes how a function’s attractiveness depends on its abundance. These parameters allow us to position diverse systems, from microorganisms to metropolitan areas, within a two-dimensional abstract space. This mapping suggests underlying commonalities and differences in the foundational mechanisms that drive the growth of these systems.
1 Introduction
@@
Fundamental to complex adaptive systems—encompassing biological organisms, human organizations, and urban areas—is the range of functions (diversity) and their relative abundance (specialization). Diversity, or the composition of a system’s sub-components, indicates a broad range of abilities and capabilities for organisms or organizations to tackle complex challenges, hedge against risks, and thus adapt to complex environments [1, 2, 3, 4, 5, 6]. As these systems mature, develop, and expand, their function diversity grows, reflecting an increase in adaptability and resilience [7, 1, 2, 8, 3]. Remarkably consistent empirical regularities are observed in both biological and socio-economic systems, despite their vastly different environmental conditions and evolutionary trajectories. Examples include the species-area curve—the empirical regularity for how the number of plant and animal species increases with land area [2]—and the scaling of microbial richness with community size [9, Fig. 1]. A similar pattern is observed in urban areas—the range of occupations and business types is strongly associated with population [4, 5, 10].
@@
While diversity provides a wide foundation of capabilities, specialization refines these capabilities for targeted functions, concentrating on a select few functions or skills to increase performance and competitive edge [11, 12]. For example, certain flowers adapt by developing nectars tailored to their specific pollinators, universities may concentrate on certain research areas like quantum computing, and cities may become hubs for certain sectors such as financial services. Specialization is manifested in the abundance distribution of functions—whether components or resources are concentrated in a few functions or evenly spread across many. Empirically, specialization has been quantified using indices like the Gini coefficient, location quotient, and entropy measures, all of which capture information in the abundance distribution [13, 14, 15, 12]. Specialization within the diversification process provides insights into how systems efficiently allocate their resources, stand out in competitive environments, and add value to the broader system’s complexity and resistance [16].
@@
Indeed, different systems pursue distinct strategic choices, objectives, and impacts under different specific scopes and subject matters. Accordingly, studies of diversity and specialization have predominantly been siloed within isolated disciplinary boundaries, thus limiting the possibility of a comprehensive understanding. Research efforts have often focused on specifics, ranging from the effects of team diversity [1, 17] to the ramifications of organizational diversity [18, 19, 20], labor specialization [21, 22], cellular heterogeneity [23, 24], or economic diversity and complexity [25, 26]. While these studies are invaluable within their respective fields, findings are often limited to a particular organizational type or the impact of single variables. This segmented approach runs the risk of missing out on identifying overarching principles and theories that underlie the processes of diversification and specialization across a variety of systems.
Thus there is a need for a comprehensive framework to study diversification and specialization across complex systems. Evidence supports broad regularities across a wide range of systems. Previous studies have identified universal patterns of abundance fluctuations across microbial communities, tropical forests, and urban populations [27]. Additionally, Heaps’ Law, which describes regularities in the scaling of novel elements, has been observed in various human-generated artifacts, including books and musical compositions, and biological components [28, 29].
The existing literature in both economics [30, 31] and biology [32, 33, 34, 35] often presents diversification and specialization as opposing poles of a linear spectrum. While this perspective is valid, especially in the context of finite resource allocation, it does not universally hold. For instance, research has demonstrated that introducing more minor entities into a market can, paradoxically, reinforce the dominance of the major players [36]. This indicates that diversification and specialization may be more accurately understood as a two-dimensional problem.
@@
In this study, we investigate function diversity and specialization in a broad range of socio-economic and biological systems, including bacteria, companies, government agencies, universities, and cities. We explore the interconnected nature of diversification and specialization across a wide range of systems by asking: Do various biological and social systems display common patterns of function diversity and abundance, and to what extent? What general governing principles could determine the function diversity and specialization of complex adaptive systems? We first measure the range (diversity) and concentration of functions in these systems (abundance distributions) within each system. We identify functional forms for each attribute to compare disparate systems. We then propose a general model for the growth of function diversity and abundance, explaining both the regularities and differences observed in the empirical data.
@@
2 Results
2.1 Empirical Results
Figure 1: The number of distinct functions versus system size (top row) and rank-frequency distribution of function abundance (bottom row) in several complex adaptive systems. (A,D) Bacterial and archaea cells, (B,E) US federal agencies, and (C,F) US metropolitan statistical areas (MSA). The dependence of the number of functions on size measure $N$ in (A) and (B) can be approximated by the power law, $D\sim N^{\beta}$. Also shown is the best-fit scaling exponent, $\beta$, and its $95\%$ confidence interval in the square brackets. (C) In contrast, logarithmic scaling occurs for MSA’s, $D\sim\log N$.
@@ 2.1 Empirical Results
Size measure Function diversity measure $\beta$ 95% CI R-sq
Bacteria Total expressed protein Distinct expressed proteins 0.35 [0.31, 0.39] 0.86
Norwegian companies Number of employees Distinct occupations 0.46 [0.45, 0.46] 0.88
US federal agencies Number of employees Distinct occupations 0.50 [0.46, 0.53] 0.87
@@
US universities, associate level Number of faculty Distinct academic programs 0.52 [0.50, 0.55] 0.61
@@
US universities, bachelor level & above Number of faculty Distinct academic programs 0.57 [0.55, 0.59] 0.75
Cities Working population Distinct occupations Logarithm scaling 0.94
@@
Our study analyzes empirical data from microorganisms to firms to metropolitan areas, including 47 bacteria and archaea cell observations [39], 125 US federal government agencies, 3,191 Norwegian companies, 2,397 US universities, and 422 US metropolitan areas. For each system, we determine the diversity, the number of distinct functions $D(N)$ and the distribution of these functions’ abundance $f_{i}$ as a function of system size $N$.
@@
We define system size as the total number of individuals within a system. In social systems, this is simply the number of people in the system, with functions represented by occupations, which are categorized using standardized classifications (see SI for details on data and methods). In bacteria and archaea, key functional components responsible for metabolism and cellular processes are proteins, analogous to workers in an organization. Accordingly, we measure system size in these biological systems as the total number of expressed proteins, while function is quantified as the number of distinct expressed proteins. Figure 1 illustrates how the range of distinct functions, $D(N)$, scales with the system size $N$. This scaling framework allows for direct comparisons across diverse systems, as demonstrated in previous studies utilizing scaling analysis [40, 41, 42]. The data for bacteria and archaea are visualized in panel A, federal agencies in panel B, and metropolitan areas in panel C. Analysis of additional datasets, including Norwegian companies and universities, are summarized in Table 1.
@@
For most systems, this scaling behavior is well-captured by a power-law relation $D=D_{0}N^{\beta}$, where $\beta$ is the scaling exponent, and $D_{0}$ is a constant. For bacteria and archaea, companies, federal agencies, universities of associate level, bachelor level and above, the scaling exponents are sublinear, ranging between $0.35$ to $0.57$. Notably, such scaling behavior resembles Heaps’ Law, which is observed in the growth of unique types across diverse human artifacts and biological components [28, 29]. Across bacteria and archaea, the consistency of this pattern is striking, given that each of the proteomes comes from organisms that have followed unique evolutionary trajectories, have different metabolisms, and occupy unique ecological niches. Comparative analyses of systems ranging from books and musical compositions to genomes, using appropriate diversity metrics, have consistently identified scaling exponents between 0.35 and 0.60, showcasing a consistent scaling pattern across both cultural artifacts and biological systems [43, 44, 45, 46, 47].
@@
Among the systems analyzed, a notable exception is metropolitan areas. Unlike other datasets, the scaling of function diversity in metropolitan areas does not follow a power-law pattern but is better captured by a logarithmic relationship, $D=b\log N+c$ , as shown in Figure 1(C). This observation suggests that different mechanisms govern the growth of function diversity in cities compared to other systems. However, the sub-linear growth of all of these systems implies that the rate of increase of their function diversity systematically slows down as the system grows in size. For example, this pattern indicates that function diversity grows more rapidly with population increase in smaller cities but decelerates as cities become larger [4].
@@
This logarithm scaling behavior carries two important implications for urban areas. First, if a city’s population grows exponentially with time ($t$), i.e., $N(t)=N_{0}e^{\delta t}$, where $N_{0}$ is the population at time $t=0$, and $\delta$ is the growth rate, then function diversity is predicted to grow linearly in time, with $D=\delta bt+D_{0}$, where $b$ is the parameter fitted from $D=b\log(N)+c$, and $D_{0}=b\log(N_{0})+c$ is a constant determined by the initial condition of the system at $t=0$. Then the growth rate of function diversity ($dD/dt$) is given by the product, $\delta\,b$, which is proportional to the population growth rate. For example, the growth rate $\delta$ for US urban areas on average is $0.71\%$ in 2019 [48] while the fitted parameter $b$ from data is $128$, which together predict that the growth rate for function diversity ($dD/dt$) is $\delta\,b=0.9$ occupations per year. Second, since most socio-economic metrics in urban areas, such as GDP, wages, and patent production, follow an approximate superlinear power-law relationship with population size [41], our findings suggest that function diversity serves as a key driver of recombinant growth [49, 50]. As functional diversity expands, the combinatorial possibilities for innovation and productivity multiply, contributing to exponential increases in economic output. This underscores the central role of diversity in sustaining and accelerating economic performance in urban systems.
@@
Lastly, we analyze the distribution of function abundance within each organization. Figure 1(D), (E), and (F) illustrate the variation in the relative abundance of functions as a function of rank for bacteria and archaea, federal agencies, and metropolitan areas, respectively. Grey dots represent all data points, while select organizations are highlighted in color for clarity.
@@
Our analysis reveals both commonalities and differences across systems. The commonality is manifested in the fact that within a given system (bacteria, companies, etc.), the range of function diversity generally grows as a sublinear power law with system size. This commonality suggests that while varying in goals, environment, and history, different organisms may share a similar underlying dynamic in how their function range is generated and grows with size. The differences across systems, however, are also important and manifested in two ways. First, the function range in cities scales differently from the other systems, suggesting the mechanisms that lead to diversity in cities may be fundamentally different from the other entities. Second, among the systems that exhibit power law scaling, there are variations in the mechanisms that give rise to the different exponents. Our analyses are the first to attempt to quantitative assessment for comparison across disparate systems.
@@
2.2 Mathematical Model
Our model aims to explain the growth dynamics of a system as new individuals join. Within this framework, a new member that joins the system can lead to one of two scenarios: (i) the new member creates a new function (diversification) or (ii) the newcomer assimilates into a pre-existing function (specialization). For example, the individuals entering the system are: newly synthesized proteins in bacteria; newly hired staff in federal agencies; and new professionals entering the job market in cities. These dynamics are aligned with the foundational principles of the Yule-Simon model [52, 53]. We mathematically represent these dynamics by postulating the following growth rule for $n_{k}$, the number of functions with $k$ individuals:
$\frac{\partial n_{k}}{\partial N}=p\delta_{k,1}+(1-p)\left(q_{k-1}n_{k-1}-q_{k}n_{k}\right)\;.$ (1)
Here $p$ is the probability of a new function being created when a new individual joins the system, $q_{k}$ denotes the probability that the new individual joins a function that currently contains $k$ individuals, and $\delta_{k,1}$ is the Dirac delta function. The first term of the equation accounts for the creation of a new function, in which the number of functions with $k=1$ individuals increases by 1. The second term accounts for an individual that joins an existing function. If the individual joins a function with $k-1$ existing individuals, $n_{k}$ increases by 1. If the individual joins a function with $k$ individuals, $n_{k}$ decreases by 1 (as that function now has $k+1$ individuals). In other scenarios, $n_{k}$ remains unchanged. Figure 2 visually summarizes this dynamic process.
Figure 2: A conceptual diagram illustrating the structure of the mathematical model. Functions are characterized by the number of individuals in them. When a new function is created, it adds to functions with 1 individual (with probability $p$ in each time step). Functions with $k$ individuals move to the next category as an individual joins this function (with probability $q_{k}$ in each time step).
@@ 2.2 Mathematical Model
2.3 Model predictions and categorization of systems by diversification and specialization parameters
We simulate a system’s growth based on Eq. 1 and compare the model’s predictions with empirical data. We then estimate the parameters $\gamma$ and $\theta$ for each system by comparing the function range and abundance generated by simulation with those from data on bacteria, federal agencies, and cities. This estimation is achieved by minimizing the Euclidean distance between normalized rank-frequency distributions in logarithmic-space using the adaptive differential evolution algorithm for optimization [59] (see SI for methods).
Figure 3 (a-c) shows the model’s predicted rank-frequency distributions are well aligned with three selected systems: Bartonella henselae (bacterial cell), the Army (federal agency), and Warren-Troy-Farmington Hills (urban area in Michigan). See SI for more examples of rank-frequency distribution predictions. The bottom row of Fig. 3 show their respective diversity relationship with those predicted by the model. Both results demonstrate good agreement across all three types of systems for both the rank-frequency distributions and the diversity scaling.
Figure 4 summarizes our simulation results, mapping over a hundred cases onto the parameter space of $\theta$ and $\gamma$. Systems cluster by type, indicating distinct within-class dynamics. At the same time, patterns emerge across systems along both $\theta$ and $\gamma$, revealing shared mechanisms that operate both within and across system classes. For instance, bacteria, federal agencies, and cities exhibit parallel trends in the specialization parameter $\gamma$, which captures how the abundance of existing functions shapes subsequent growth. All three systems consistently show positive $\gamma$ values, but predominantly sublinear (below 1), indicating a general tendency toward diminishing returns in specialization—growth is less likely to concentrate solely on already dominant functions. Nevertheless, the variation in $\gamma$ is wider for federal agencies and bacteria than for cities, suggesting that urban systems may exhibit a more uniform—and possibly universal—specialization pattern. On the other hand, $\theta$ has a broader range for different class of systems than $\gamma$.

### v3

@@ Abstract
Function diversity, the range of tasks individuals perform, and specialization, the distribution of function abundances, are fundamental to complex adaptive systems. In the absence of overarching principles, these properties have appeared domain-specific. Here, we introduce an empirical framework and a mathematical model for the diversification and specialization of functions across disparate systems, including bacteria, federal agencies, universities, corporations, and cities. We find that the number of functions grows sublinearly with system size, with exponents from 0.35 to 0.57, consistent with Heaps’ Law. In contrast, cities exhibit logarithmic scaling. To explain these empirical findings, we generalize the Yule-Simon model by introducing two key parameters: a diversification parameter that characterizes how existing functions inhibit the creation of new ones, and a specialization parameter that describes how a function’s attractiveness depends on its abundance. Our model enables cross-system comparisons, from microorganisms to metropolitan areas. The analysis suggests that what drives the creation of new functions depends on the system’s goals and structure: federal agencies tend to ensure comprehensive coverage of necessary functions; cities tend to slow the creation of new occupations as existing ones expand; and cells occupy an intermediate position. Once functions are introduced, their growth follows a remarkably universal pattern across all systems.
Significance Statement
Diversification and specialization are central to complex adaptive systems, yet overarching principles across domains remain elusive. We introduce a general theory that unifies diversity and specialization across disparate systems, including microbes, federal agencies, companies, universities, and cities, characterized by two key parameters. We show from extensive data that function diversity scales with system size as a sublinear power law—resembling Heaps’ law—in all but cities, where it is logarithmic. Our theory explains both behaviors and suggests that function creation depends on system goals and structure: federal agencies tend to ensure functional coverage; cities slow new function growth as old ones expand, and cells occupy an intermediate position. Once functions are introduced, their growth follows a remarkably universal pattern across all systems.
Introduction
@@
Fundamental to complex adaptive systems—encompassing biological organisms, human organizations, and urban areas—is the range of functions (diversity) and their relative abundance (specialization). Diversity, the number of unique sub-components within a system, indicates a broad range of abilities and capabilities for organisms or organizations to tackle complex challenges, hedge against risks, and thus adapt to complex environments (see Table 1 for some examples)[57, 79, 64, 81, 48, 41]. As these systems mature, develop, and expand, their function diversity grows, reflecting an increase in adaptability and resilience [71, 57, 79, 42, 64]. Remarkably consistent empirical regularities are observed in both biological and socio-economic systems, despite their distinct environmental conditions and evolutionary trajectories. Examples include the species-area curve—the empirical regularity for how the number of plant and animal species increases with land area [79]—and the scaling of microbial richness with community size [46]. A similar pattern is observed in urban areas—the range of occupations and business types is strongly associated with population [81, 48, 33].
@@
While diversity provides a wide foundation of unique capabilities, specialization refines these capabilities for targeted functions, concentrating on a select few functions or skills to increase performance and competitive edge [62, 40]. For example, certain flowers adapt by developing nectars tailored to their specific pollinators, universities may concentrate on specific research areas like quantum computing, and cities may become hubs for certain sectors such as financial services. Specialization is manifested in the abundance distribution of functions—whether components or resources are concentrated in a few functions or evenly spread across many. Empirically, specialization has been quantified using indices like the Gini coefficient, location quotient, and entropy measures, all of which capture information in the abundance distribution [23, 29, 59, 40]. Specialization within the diversification process provides insights into how systems efficiently allocate resources and stand out in competitive environments [58].
@@
Indeed, different systems pursue distinct strategic choices, objectives, and impacts under different specific scopes and subject matters. Accordingly, studies of diversity and specialization have predominantly been siloed within isolated disciplinary boundaries, thus limiting the possibility of a comprehensive understanding. The existing literature in both economics [9, 38] and biology [50, 19, 51, 20] often presents diversification and specialization as opposing poles of a linear spectrum. While this perspective is valid, especially in the context of finite resource allocation, it does not universally hold. For instance, research has demonstrated that introducing more minor competitors into a market can, paradoxically, reinforce the dominance of the major players [17]. This indicates that diversification and specialization may be more accurately understood as a two-dimensional problem. In addition, research efforts have often focused on specifics, ranging from the effects of team diversity [57, 54] to the ramifications of organizational diversity [60, 35, 53], labor specialization [10, 36], cellular heterogeneity [22, 31, 2], or economic diversity and complexity [27, 30]. While these studies are invaluable within their respective fields, findings are often limited to a particular organizational type or the impact of single variables. This segmented approach runs the risk of missing out on identifying overarching principles and theories that underlie the processes of diversification and specialization across a variety of systems.
Thus, there is a need for a comprehensive framework to study diversification and specialization across complex systems. Classically there have been two separate approaches to characterizing commonalities in complex systems: the distributional approach [8, 16] and the feature scaling approach [28, 78, 77, 12]. The first tries to connect generative mechanisms to the shape of abundance and probability distributions, the second connects mechanisms to how aggregated properties scale with the overall size of a system often in the form of a power law $Y=Y_{0}N^{\beta}$, where $Y$ is an aggregated quantity, $N$ the system size, and $Y_{0}$ a normalization constant. Power-law scaling of aggregated properties is widely argued to be the product of adaptive systems evolving under fixed constraints that yield particular optima [78, 77, 12]. Distributional regularities are argued to stem from many processes including self-organized criticality [5, 6], simple growth and attachment rules leading to power-law degree distributions in networks [8], or sampling processes in, for example, ecological systems [37, 26].
Urban systems provide a rich example of how feature scaling relationships have proven helpful for building scientific understanding of complex, highly variable systems. Systematic scaling laws linking socioeconomic outputs and infrastructure costs to population size have inspired models proposing mechanisms including social network effects [14, 80], geometric and spatial constraints [55], and mobility patterns [66]. Although the scaling exponents of different urban quantities can vary, such differences can arise from common generative principles [80]. Studies have also examined diversity in urban systems using this framework. While new economic functions often originate in large urban centers and subsequently diffuse to smaller ones [63, 32], the diversity of business establishments and occupations exhibits striking regularities across sizes [81, 13].
Beyond urban areas, scaling laws have also been used for cross-system studies. A well-studied regularity is Heaps’ law, an empirical pattern originally observed in linguistics in how the number of unique words in a corpus scales with its length, showing a sublinear power law [28]. Similar scaling relationships were later observed in analyses from books in various languages to musical compositions to genomes—using appropriate diversity metrics, they have consistently identified power-law scaling for how number of unique elements scale with size, with exponents between 0.35 and 0.60 [52, 74, 11, 72, 61, 69]. The robustness of such empirical scaling laws in novelties within human artifacts has motivated theoretical efforts to uncover their generative mechanisms. Heaps’ law is closely related to abundance distributions, as it can be formally derived from Zipf’s law [75]. Two main classes of explanations have been proposed—one emphasizing how existing elements open adjacent possibles that enable novel elements to emerge [74, 47], and another focusing on how early commitments constrain subsequent possibilities, limiting what can be added later [18]. Together, these findings suggest that scaling regularities may reflect universal principles governing how complex systems diversify and specialize as they grow. This conceptual foundation motivates our empirical and modeling analysis of how functional diversity and specialization scale with system size across biological and socioeconomic domains.
@@
In this study, we investigate function diversity and specialization in a broad range of socio-economic and biological systems, including bacteria, companies, government agencies, universities, and cities. We explore the interconnected nature of diversification and specialization across a wide range of systems by asking: Do various biological and social systems display common patterns of function diversity and abundance, and to what extent? What general governing principles could determine the function diversity and specialization of complex adaptive systems? We first measure the range (diversity) and concentration of functions (abundance distributions) within each system. We identify functional forms for each attribute to compare disparate systems. We then propose a general model for the growth of function diversity and abundance, explaining both the regularities and differences observed in the empirical data.
@@
Results
Empirical Results
Figure 1: The number of distinct functions versus system size (top row) and rank-frequency distribution of function abundance (bottom row) in several complex adaptive systems. (A,D) Bacterial and archaea cells, (B,E) US federal agencies, and (C,F) US metropolitan statistical areas (MSA). The dependence of the number of functions on size measure NN in (A) and (B) can be approximated by the power law, $D\sim N^{\beta}$. Also shown is the best-fit scaling exponent, $\beta$, and its $95\%$ confidence interval in the square brackets. (C) In contrast, logarithmic scaling occurs for MSA’s, $D\sim\log N$. In (D–F), grey dots represent all available data points, while select entities are highlighted in color.
@@ Results
Size measure Function diversity measure $\beta$ 95% CI R-sq Obs
Bacteria & archaea Total expressed protein Distinct expressed proteins 0.35 [0.31, 0.39] 0.86 47
Norwegian companies Number of employees Distinct occupations 0.46 [0.45, 0.46] 0.88 3,191
US federal agencies Number of employees Distinct occupations 0.50 [0.46, 0.53] 0.87 125
@@
US universities, associate level Number of faculty Distinct academic programs 0.52 [0.50, 0.55] 0.61 1,069
@@
US universities, bachelor level & above Number of faculty Distinct academic programs 0.57 [0.55, 0.59] 0.75 1,328
Cities Working population Distinct occupations Logarithmic scaling 0.94 422
We define a function as the kind of task an individual component within a system is capable of performing, and function diversity as the number of distinct task types present in the system. This definition is intentionally agglomerated to enable comparisons across structurally and mechanistically diverse systems, ranging from cells to cities. While the internal architectures and generative mechanisms of these systems differ—such as gene regulation in cells, hiring decisions in firms, and curriculum development in universities—we focus on their realized functional repertoires: the observed set of distinct roles or activities that contribute to system-level capabilities.
In biology, proteins are often referred to as the workhorses of the cell because they carry out nearly all of its essential functions. Each type of protein performs a specific task. For example, enzymes catalyze biochemical reactions to produce particular metabolites; motor proteins move substances within the cell; transcription factors are regulatory proteins that control when other proteins are expressed, preventing the simultaneous production of proteins that could interfere with one another. These protein roles have analogs in many socioeconomic systems, such as factory production workers, delivery workers, and managers. Occupations in cities and firms capture the range of economic activities their components can carry out, while academic programs in universities reflect the range of instructional or research activities the institution can perform. Across domains, this functional repertoire reflects the system’s capacity to adapt, specialize, and perform complex tasks. By analyzing how function diversity scales with system size, we seek to uncover shared statistical regularities that may point to general principles.
@@
Our study analyzes empirical data from microorganisms to firms to metropolitan areas, including 47 bacteria and archaea cell observations, 125 US federal government agencies, 3,191 observations of Norwegian companies, 2,397 US universities, and 422 US metropolitan areas. For each system, we determine the diversity, the number of distinct functions $D(N)$ and the distribution of these functions’ abundance as a function of system size $N$.
@@
We define system size as the total number of individuals within a system. In social systems, this is simply the number of people in the system, with functions represented by occupations, which are categorized using standardized classifications (see SI for details on data and methods). In bacteria and archaea, key functional components responsible for metabolism and cellular processes are proteins, analogous to workers in an organization. Accordingly, we measure system size in these biological systems as the total number of expressed proteins, while function is quantified as the number of distinct expressed proteins. Figure 1 illustrates how the range of distinct functions, $D(N)$, scales with the system size $N$. This scaling framework allows for direct comparisons across diverse systems, as demonstrated in previous studies utilizing scaling analysis [78, 12, 77]. The data for bacteria and archaea are visualized in panel A, federal agencies in panel B, and metropolitan areas in panel C. Analysis of additional datasets, including Norwegian companies and universities, are summarized in Table 1, with visualization of scaling relationships included in the SI.
@@
For most systems, this scaling behavior is well-captured by a power-law relation $D=D_{0}N^{\beta}$, where $\beta$ is the scaling exponent, and $D_{0}$ is a constant. For bacteria and archaea, companies, federal agencies, and universities (both associate and bachelor level and above), the scaling exponents are sublinear, ranging between $0.35$ to $0.57$. This consistency is particularly striking in bacteria and archaea, given that each of the proteomes comes from organisms that have followed unique evolutionary trajectories, have different metabolisms, and occupy unique ecological niches. The sublinear nature of this scaling ($\beta<1$) further implies that as systems grow in size, the increase in the number of distinct functions is less than proportional. This suggests a tendency for larger systems to reuse or replicate existing functions rather than continuously introducing new ones.
Notably, this behavior closely resembles Heaps’ Law, the sublinear power-law scaling for how the number of unique elements scale with size in entities ranging from books in various languages to musical compositions to genomes. While Heaps’ Law in human systems is usually associated with the accumulation of novelty in produced artifacts—such as vocabulary in texts or motifs in music—our findings suggest that this same empirical regularity extends to the structure of many human organizations themselves. In other words, the sublinear growth of diversity is not limited to what humans create, but also characterizes many agglomerations they form, such as companies and universities. This shift—from artifacts to collectives—may suggest a deeper principle underlying how complexity scales in systems shaped by human activity.
@@
Among the systems analyzed, a notable exception is metropolitan areas. Unlike other datasets, the scaling of function diversity in metropolitan areas does not follow a power-law pattern but is better captured by a logarithmic relationship, $D=b\log N+c$ , as shown in Figure 1(C) (in SI, we show this logarithmic scaling relationship is robust over time). This observation suggests that different mechanisms govern the growth of function diversity in metropolitan areas compared to other systems. Function diversity grows more rapidly with population in smaller cities but decelerates as cities become larger. Notably, the logarithmic scaling observed in metropolitan areas departs from Heaps’ Law. This deviation indicates that Heaps-like behavior is not ubiquitous, but instead contingent on structural factors that shape how systems diversify and specialize. If a fundamental underlying mechanism for diversity and specialization exists, it must be capable of explaining both scaling behaviors—under some conditions, Heaps-like behavior emerges, while in others, diversification and specialization follow alternative patterns.
@@
This logarithmic scaling behavior carries two important implications for urban areas. First, if a city’s population grows exponentially with time ($t$), i.e., $N(t)=N_{0}e^{\alpha t}$, where $N_{0}$ is the population at time $t=0$, and $\alpha$ is the growth rate, then function diversity is predicted to grow linearly in time, with $D=\alpha bt+D_{0}$, where $b$ is the parameter fitted from $D=b\log(N)+c$, and $D_{0}=b\log(N_{0})+c$ is a constant determined by the initial condition of the system. Then the growth rate of function diversity ($dD/dt$) is given by the product, $\alpha\,b$, which is proportional to the population growth rate. For example, the growth rate $\alpha$ for US urban areas on average is $0.71\%$ in 2019 [73] while the fitted parameter $b$ from data is $128$, which together predict that the growth rate for function diversity ($dD/dt$) is $\alpha\,b=0.9$ occupations per year. Second, since most socio-economic metrics in urban areas, such as GDP, wages, and patent production, follow an approximate superlinear power-law relationship with population size [12], our findings suggest that function diversity serves as a key driver of recombinant growth [76, 82]. As functional diversity expands, the combinatorial possibilities for innovation and productivity multiply, contributing to exponential increases in economic output. This underscores the central role of diversity in sustaining and accelerating economic performance in urban systems.
@@
Lastly, we analyze the distribution of function abundance within each organization. Figure 1(D), (E), and (F) illustrate the variation in the relative abundance of functions as a function of rank for bacteria and archaea, federal agencies, and metropolitan areas, respectively. Grey dots represent all data points, while select entities are highlighted in color for clarity.
@@
Our analysis reveals both commonalities and differences across systems. The commonality is manifested in the fact that within a given system (bacteria, companies, etc.), the range of function diversity generally grows as a sublinear power law with system size. This commonality suggests that while varying in goals, environment, and history, different organisms may share a similar underlying dynamic in how the number of unique functions grows with size. The differences across systems, however, are also important and manifested in two ways. First, the number of unique functions in cities scales differently from the other systems, suggesting the mechanisms that lead to diversity in cities may be fundamentally different from the other entities. Second, among the systems that exhibit power-law scaling, there are variations in the scaling exponents. Our analyses provide the first quantitative assessment enabling comparison across disparate systems.
@@
Mathematical Model
Our model extends the foundational principles of the Yule-Simon model [83, 70], to describe the growth dynamics of a system as new individuals join. Within this framework, a new member that joins the system can lead to one of two scenarios: (i) the new member creates a new function (diversification, with probability denoted as $p$) or (ii) the newcomer assimilates into a pre-existing function (specialization, with probability 1p1-p). For example, the individuals entering the system are: newly synthesized proteins in a bacteria; newly hired staff in a federal agency; and a new professional entering the labor market of a city. We assume Markovian dynamics throughout the model, which implies that the probability of adding a new function, or expanding an existing function, is only dependent on the present state of the abundance distributions. Figure 2 visually summarizes this dynamic process.
Figure 2: Conceptual diagram of the mathematical model. As each new individual joins the system, it either creates a new function with probability $p$, where $p$ depends on the existing abundance distribution; or joins an existing function with probability 1p1-p. When it joins an existing function, the probability of joining a particular existing function with $k$ members, $q_{k}$, follows a generalized preferential attachment process.
@@
Model predictions and categorization of systems by diversification and specialization parameters
We simulate a system’s growth based on Eqs. 1 and 2 and compare the model’s predictions with empirical data. The dynamics of the model are shown in Figure 2. Systems of various sizes are treated as points along the model’s growth trajectory, representing successive stages in the accumulation of individuals. This assumes that larger systems arise through the sequential addition of individuals to smaller ones. Under this framework, the dynamic model generates cross-sectional predictions that can be directly compared to observed patterns of diversity and abundance across systems of varying sizes.
We then estimate the parameters $\gamma$ and $\theta$ for each system by comparing the function range and abundance generated by simulation with those from data on bacteria, federal agencies, and cities. This estimation is achieved by minimizing the Euclidean distance between normalized rank-frequency distributions in logarithmic-space using the adaptive differential evolution algorithm for optimization (see SI for methods).
Figure 3 (a-c) shows the model’s predicted rank-frequency distributions are well aligned with three selected systems: Bartonella henselae (bacterial cell), the Army (federal agency), and Warren-Troy-Farmington Hills (urban area in Michigan). See SI for more examples of rank-frequency distribution predictions. The bottom row of Figure 3 show their respective diversity relationship with those predicted by the model. Both results demonstrate good agreement across all three types of systems for both the rank-frequency distributions and the diversity scaling.
Figure 4 summarizes our calibration results, mapping over a hundred cases onto the parameter space of $\theta$ and $\gamma$, where each marker represents the calibrated values of one entity, and marker shapes distinguish system type. Systems cluster by type, indicating distinct within-class behavior. The clustering patterns in Figure 4 reveal systematic differences in how functional specialization and diversification unfold across system types. Cities occupy a particularly narrow region of parameter space, centered around $\gamma\approx 0.8$ and $\theta\approx 1$, suggesting a highly consistent balance between diversification suppression and sublinear growth of existing functions. This tight grouping is likely associated with the uniformity in cities’ abundance distributions compared to other entities (Fig. 1(F)). In contrast, bacteria and federal agencies span a much broader region of the parameter space, likely reflecting differences in internal constraints and evolutionary or administrative heterogeneity. Together, these results indicate that cities are not only distinct but also unusually consistent in their system-level patterns compared with other entities.